discuss the stages of chronic kidney disease. What are causes and symptoms associated with chronic kidney disease? Additionally, please identify methods of monitoring and treating patients with chronic kidney disease. please provide references
Answer #1

Ans) Stages of Chronic Kidney disease:

Stage 1 with normal or high GFR (GFR > 90 mL/min)
Stage 2 Mild CKD (GFR = 60-89 mL/min)
Stage 3A Moderate CKD (GFR = 45-59 mL/min)
Stage 3B Moderate CKD (GFR = 30-44 mL/min)
Stage 4 Severe CKD (GFR = 15-29 mL/min)
Stage 5 End Stage CKD (GFR <15 mL/min)

- In the United States the two leading causes of kidney failure, also called end stage kidney disease or ESRD, are diabetes (also called Type 2, or adult onset diabetes) and high blood pressure. When these two diseases are controlled by treatment, the associated kidney disease can often be prevented or slowed down.

- Common symptoms include blood in urine, high blood pressure, and fatigue. Causes include diabetes and specific kidney diseases, which includes polycystic kidney disease. There is no cure for chronic kidney disease, which means treatment is focussed on reducing symptoms.

- Treatments may include: High blood pressure medications. Medications to lower cholesterol levels. Medications to treat anemia. Medications to relieve swelling. Medications to protect your bones. A lower protein diet to minimize waste products in your blood.